# TexSenseGAN: A User-Guided System for Optimizing Texture-Related Vibrotactile Feedback Using Generative Adversarial Network

### v4

@@ V-C User Study
However, spectrograms alone cannot fully capture the subjective experiences of the subjects. Therefore, we mainly focused on analyzing the results of Stage II, the classification task. From Fig. 12, we can observe an overview of the result of the classification task. In the five-category task, subjects achieved an accuracy of 55.9% when judging the generated vibrations (experimental group, experiment 2), compared to approximately 72.8% for the real vibrations (control group, experiment 1). The control group results indicate that subjects can distinguish between different surface vibration stimuli. Fig. 12 shows that subjects in the generation group and the control group exhibited similar accuracy trends for these 5 types of different vibrations. This preliminary evidence suggests that our TexSenseGAN can generate distinct vibration patterns that subjects can differentiate.
@@
Additionally, we can draw some conclusions based on each subject’s performance in both the generated and control group, as shown in Fig. 15. We calculated the average accuracy of each subject for each class, obtained 100 sets of accuracy. The figure roughly shows an interesting trend: subjects who were good at distinguishing real vibrations in the control group also tended to have higher classification accuracy in the generated group. To further illustrate this phenomenon, we constructed a multivariate regression model following the process below to verify the relationship between the accuracy in the control group and the accuracy in the generated group.
Figure 13: The confusion matrix of the classification task of generated vibrations.
Figure 14: The confusion matrix of the classification task of real vibrations.
@@ VI Discussion
Therefore, in future work, we plan to explore other generative models to enhance the diversity and precision of the generated vibrations. For example, incorporating diffusion models can help improve the variety and fidelity of the output [37]. Diffusion models can be easy to train compared with GAN and can provide high quality results. Additionally, the current optimization process is time consuming and requires subjects to focus intensely on detecting subtle differences while adjusting the slider position. Streamlining the interface for more intuitive operation can significantly improve user experience. Our current presentation system is designed specifically to deliver texture-related vibratory feedback on simple devices, such as game controllers, rather than to replicate the sensation of real textures themselves. Therefore, in future research, we may further explore the application of this framework to spatial and temporal stimuli and aim to deploy it on equipment better suited to the original data. For instance, we envision using devices like the Phantom of 3D Systems, to fit the data collection method better and consider spatial stimuli to create more realistic tactile feedback.

### v5

@@ V-C Subjective Similarity and Generation Results
However, spectrograms alone cannot fully capture the subjective experiences of the subjects. Therefore, we mainly focused on analyzing the results of Stage II, the classification task.
V-D Inter-class Classification
From Fig. 12, we can observe an overview of the result of the inter-class classification task. In the five-category task, subjects achieved an accuracy of 55.9% when judging the generated vibrations (experimental group, experiment 2), compared to approximately 72.8% for the real vibrations (control group, experiment 1). The control group results indicate that subjects can distinguish between different surface vibration stimuli. Fig. 12 shows that subjects in the generation group and the control group exhibited similar accuracy trends for these 5 types of different vibrations. This preliminary evidence suggests that our TexSenseGAN can generate distinct vibration patterns that subjects can differentiate.
@@ V-D Inter-class Classification
V-E Subjective Classification and Mathematical Similarity
To assess the effectiveness of subjective judgment as the primary guide for adjustment, we computed the objective similarity between generated and target samples as a supplementary measure. Since our primary concern is whether the frequency components of the samples are similar, and given that the generated and target samples are not entirely identical (e.g., differences in starting time and phase), we applied a Fourier transform to all samples and extracted only their magnitude spectra. For each generated sample, we computed cosine similarity with the magnitude spectra of all target samples. The resulting pairwise similarities were averaged to form the matrix shown in Fig. 13.
Comparing Fig. 13 with Fig. 13 and Fig. 13, the generated samples perform noticeably worse on objective metrics than on subjective classification results. However, we also observe that for some easily confused categories, the objective metrics exhibit trends similar to those seen in subjective classification, indicating partial agreement between mathematical and perceptual measures. In other words, the trend of similarity provides a common ground between subjective and mathematical indicators, reinforcing each other. This phenomenon suggests that within the framework of this study, subjective perception is more sensitive than mathematical metrics, allowing us to obtain more meaningful and precise results.
Figure 14: The average accuracy of each subject.
V-F Regression Analysis
We can draw some conclusions based on each subject’s performance in both the generated and control group, as shown in Fig. 14. We calculated the average accuracy of each subject for each class, obtained 100 sets of accuracy. The figure roughly shows an interesting trend: subjects who were good at distinguishing real vibrations in the control group also tended to have higher classification accuracy in the generated group. To further illustrate this phenomenon, we constructed a multivariate regression model following the process below to verify the relationship between the accuracy in the control group and the accuracy in the generated group.
@@ VI Discussion
Therefore, in future work, we plan to explore other generative models to enhance the diversity and precision of the generated vibrations. For example, incorporating diffusion models can help improve the variety and fidelity of the output [36]. Diffusion models can be easy to train compared with GAN and can provide high quality results. Additionally, the current optimization process is time consuming and requires subjects to focus intensely on detecting subtle differences while adjusting the slider position. Streamlining the interface for more intuitive operation can significantly improve user experience. Our current presentation system is designed specifically to deliver texture-related vibratory feedback on simple devices, such as game controllers, rather than to replicate the sensation of real textures themselves. Therefore, in future research, we may further explore the application of this framework to spatial and temporal stimuli and aim to deploy it on equipment better suited to the original data. For instance, we envision using devices like the PHANToM, to fit the data collection method better and consider spatial stimuli to create more realistic tactile feedback.